A bowling ball (solid sphere) of mass 5.0 kg and radius of 11cm. What is theangular acceleration if a force of 8.0 N is applied to the surface (edge) of theball?
Answer #1

mass of the sphere, m = 5.0 kg

radius, r = 11 cm =0.11 m

Moment of inertia of the solid sphere, J = (2/5)*m*r2 = (2/5)*(5.0)*(0.11)2 = 0.0242 kg.m2

Torque produced by the force applied to the surface of the ball, T = F*r = (8.0 * 0.11) = 0.88 N.m

Angular acceleration of the ball, a = ( T / J ) = ( 0.88 / 0.0242 ) =36.364 rad/s2 (Ans)